suppose you wish to store 1.0 kW.s in a parallel plate capicator when a potential difference of 10V is applied across its plates What must be the capacitance of your parallel plate capacitor


1
Expert's answer
2021-02-22T10:24:43-0500

Energy stored in a parallel plate capacitor"(U)= \frac{1}{2}CV^2"

Given, "U= 1.0\ KW.s=1000\ J"

"V= 10\ V"

Substituting these values in above equation,

"1000=\frac{1}{2}C(10)^2\implies C=20\ F"